# Purely gravitational dark matter production in warm inflation

### v1

@@ Appendix A Derivation of the Boltzmann equation
×$\times\left(2\pi\right)^{4}\delta^{4}\left(p_{1}+p_{2}-p_{3}-p_{4}\right)\sum_{\mathrm{pol}}\left|\mathcal{M}\right|^{2}$
×$\times\left[f_{1}f_{2}\left(1\pm f_{3}\right)\left(1\pm f_{4}\right)-f_{3}f_{4}\left(1\pm f_{1}\right)\left(1\pm f_{2}\right)\right],$ (59)
@@
is the distribution function, where $+1$ corresponds to fermion and $-1$ corresponds to boson. For the process we consider, the chemical potential $\mu_{1,2}\gg\mu_{3,4}$, and the main contribution comes from particles with $E-\mu\gg T$, therefore $f\ll 1$ and $f_{1,2}\gg f_{3,4}$. Then Eq. (A) can be simplified as:
@@
×$\times\left(2\pi\right)^{4}\delta^{4}\left(p_{1}+p_{2}-p_{3}-p_{4}\right)\sum_{\mathrm{pol}}\left|\mathcal{M}\right|^{2}f_{1}f_{2}.$ (61)
@@
×$\times\sigma\sqrt{\left(p_{1}\cdot p_{2}\right)^{2}-m_{1}^{2}m_{2}^{2}}.$ (63)
@@
×$\times\sigma\sqrt{\left(p_{1}\cdot p_{2}\right)^{2}-m_{i}^{4}}.$
[…]
×$\times\sigma\left(s-4m_{i}^{2}\right)\sqrt{E_{+}^{2}-s}$
[…]

### v3

@@ Appendix A Derivation of the Boltzmann equation
$\times\left(2\pi\right)^{4}\delta^{4}\left(p_{1}+p_{2}-p_{3}-p_{4}\right)\sum_{\mathrm{pol}}\left|\mathcal{M}\right|^{2}$
$\times\left[f_{1}f_{2}\left(1\pm f_{3}\right)\left(1\pm f_{4}\right)-f_{3}f_{4}\left(1\pm f_{1}\right)\left(1\pm f_{2}\right)\right],$ (61)
@@
is the distribution function, where +1+1 corresponds to fermion and 1-1 corresponds to boson. For the process we consider, the chemical potential $\mu_{1,2}\gg\mu_{3,4}$, and the main contribution comes from particles with $E-\mu\gg T$, therefore $f\ll 1$ and $f_{1,2}\gg f_{3,4}$. Then Eq. (A) can be simplified as:
@@
$\times\left(2\pi\right)^{4}\delta^{4}\left(p_{1}+p_{2}-p_{3}-p_{4}\right)\sum_{\mathrm{pol}}\left|\mathcal{M}\right|^{2}f_{1}f_{2}.$ (63)
@@
$\times\sigma\sqrt{\left(p_{1}\cdot p_{2}\right)^{2}-m_{1}^{2}m_{2}^{2}}.$ (65)
@@
$\times\sigma\sqrt{\left(p_{1}\cdot p_{2}\right)^{2}-m_{i}^{4}}.$
$=$~ $\frac{g_{i}^{2}}{32\pi^{4}}\int\mathrm{d}s\int\mathrm{d}E_{+}e^{-E_{+}/T}$
$\times\sigma\left(s-4m_{i}^{2}\right)\sqrt{E_{+}^{2}-s}$
$=$~ $\frac{Tg_{i}^{2}}{32\pi^{4}}\int\mathrm{d}s\sigma\sqrt{s}\left(s-4m_{i}^{2}\right)K_{1}\left(\frac{\sqrt{s}}{T}\right)$
$\equiv$~ $\mathcal{R}(T).$ (72)